# Uncertainty Visualization via Low-Dimensional Posterior Projections

### v1

@@ Abstract
In ill-posed inverse problems, it is commonly desirable to obtain insight into the full spectrum of plausible solutions, rather than extracting only a single reconstruction. Information about the plausible solutions and their likelihoods is encoded in the posterior distribution. However, for high-dimensional data, this distribution is challenging to visualize. In this work, we introduce a new approach for estimating and visualizing posteriors by employing energy-based models (EBMs) over low-dimensional subspaces. Specifically, we train a conditional EBM that receives an input measurement and a set of directions that span some low-dimensional subspace of solutions, and outputs the probability density function of the posterior within that space. We demonstrate the effectiveness of our method across a diverse range of datasets and image restoration problems, showcasing its strength in uncertainty quantification and visualization. As we show, our method outperforms a baseline that projects samples from a diffusion-based posterior sampler, while being orders of magnitude faster. Furthermore, it is more accurate than a baseline that assumes a Gaussian posterior.
1 Introduction
Figure 1: Informed uncertainty visualization. Point estimation methods receive a distorted image and output only a single solution, e.g., the MMSE estimator (top). NPPC [49] complements MMSE estimators with input-adaptive uncertainty directions (principal components of the posterior) without modeling output likelihood (middle). Our method (bottom) learns the input-adaptive projected posterior distribution, facilitating a likelihood-informed uncertainty visualization.
@@ 1 Introduction
Recently, several works proposed to visualize posterior uncertainty by traversing along the principal components (PCs) of the posterior [49, 4]. These visualizations shed light on the main modes of variation along which the solution could vary, given the input, and have thus been demonstrated to be natural for signals with strong correlations, like images. These methods are also mathematically appealing since one-dimensional statistics of the projected posterior along the PCs (like variances or quantiles) provide a tighter and more accurate description of the underlying uncertainty [4]. Nonetheless, methods in this category do not provide estimates for the likelihoods of the projected solutions within the low-dimensional subspace. This hinders the user’s ability to interpret and visualize the effective support of the projected posterior.
@@
In this work, we propose to model the projected posterior using a conditional energy-based model (EBM) that receives a degraded measurement and an affine subspace (parameterized by an origin point and a set of directions) and can output the projection of the posterior probability for any queried point within that subspace. This allows us to visualize the projected posterior within any low-dimensional space (e.g., the one spanned by the dominant posterior PCs), facilitating informed navigation of posterior uncertainty in both one and two dimensions (Fig. 1). Our method, which we coin projected posterior distribution estimation (PPDE), is a general technique for uncertainty visualization that is seamlessly transferrable across tasks and datasets. Here, we illustrate the method with an MSE-optimized point prediction and the posterior PCs provided by NPPC [49]. However, it is not constrained to this specific choice and can wrap around any point estimate and an accompanying set of directions.
@@
As we illustrate, our proposed visualization approach cannot be achieved with methods that explicitly learn the high-dimensional posterior distribution, like conditional invertible models (e.g., SRFlow) or score-based models that enable the calculation of the exact likelihood using the probability flow ODE [66, 67]. In particular, the naive approach of taking a 2D or 1D slice through the high-dimensional distribution gives meaningless results because it almost always misses the high-probability regions in space (see Fig. 3 and the explanation in Sec. 3.1). By contrast, our approach corresponds to projecting the distribution onto the 2D or 1D slice (i.e., integrating over all other dimensions), an effect that is practically impossible to achieve with models that output the density in the high-dimensional space.
@@
We demonstrate the practical benefit of our method on multiple inverse problems in imaging. We quantitatively compare our learned posteriors to a Gaussian approximation obtained from [49], as well as to a baseline that projects samples generated by a diffusion-based posterior sampler and applies kernel-density estimation (KDE) in the low-dimensional space. In both cases, we show a significant improvement in sample log-likelihood across tasks and datasets. This illustrates the benefit offered by our method compared to works that approximate the posterior with a Gaussian [14, 47, 44, 50] or attempt to rely on posterior samplers [4], setting forth a new approach to proper uncertainty visualization.
2 Related Work
@@
3 Method
Figure 2: Notations illustration. An image $\bm{x}$ is projected onto a measurement-adaptive affine subspace $\mathcal{A}(\bm{y})$, and represented by projection coefficients $\bm{v}=[v_{1},v_{2}]^{\top}$. The restored image within the subspace is given by $\bm{x}_{\text{restored}}$.
3.1 Projecting the posterior distribution
Our goal is to visualize the uncertainty when predicting a signal $\bm{x}$ based on measurements $\bm{y}$. In the context of imaging, $\bm{y}$ often represents a degraded version of $\bm{x}$ (e.g., noisy, blurry). We assume that $\bm{x}$ and $\bm{y}$ are realizations of random vectors $\mathbf{x}$ and $\mathbf{y}$, respectively. The desired uncertainty we aim to explore is therefore encapsulated in the posterior distribution $p_{\mathbf{x}|\mathbf{y}}(\bm{x}|\bm{y})$. As mentioned earlier, this distribution is typically defined over a high-dimensional space, and therefore visualizing it requires projection onto a lower dimensional manifold (e.g., 1D or 2D).
@@ 3.1 Projecting the posterior distribution
$\left\{\bm{x}:\bm{x}=\bm{x}_{0}(\bm{y})+\bm{W}(\bm{y})\bm{v},\;\bm{v}\in\mathbb{R}^{K}\right\},$ (1)
@@
$\bm{W}(\bm{y})=\begin{bmatrix}|&|&&|\\ \bm{w}_{1}(\bm{y})&\bm{w}_{2}(\bm{y})&\dots&\bm{w}_{K}(\bm{y})\\ |&|&&|\end{bmatrix}.$ (2)
@@
$\mathcal{A}(\bm{y})=\left\{\bm{x}_{0}(\bm{y}),W(\bm{y})\right\}$ (3)
and with slight abuse of terminology, refer to $\mathcal{A}(\bm{y})$ as our affine subspace (or just subspace). Our method can wrap around any model that outputs a subspace $\mathcal{A}(\bm{y})$ for each input $\bm{y}$. Here we use the NPPC method [49] (see Sec. 3.2).
For any given input $\bm{y}$, our goal is to visualize the posterior distribution of the coordinates of the projection of $\mathbf{x}$ onto the subspace $\mathcal{A}(\bm{y})$, which we denote by
$\mathbf{v}=\bm{W}(\bm{y})^{\top}(\mathbf{x}-\bm{x}_{0}(\bm{y})).$ (4)
@@
$\bm{x}_{\text{restored}}=\bm{x}_{0}(\bm{y})+\bm{W}(\bm{y})\bm{v}.$ (5)
Figure 2 summarizes our notations.
@@
Figure 3: Posterior slicing vs. projection. (a) Schematic illustrating the difference between the slice and the projection of a high-dimensional manifold onto a low-dimensional subspace. (b) Sliced and projected posterior comparison for a 1D affine subspace defined by the posterior mean and the first PC in the task of digit inpainting. The sliced posterior has been computed using an EBM.
@@
It is instructive to note the difference between the PDD and the sliced posterior distribution (i.e., the posterior evaluated on $\mathcal{A}(\bm{y})$). Any $\bm{x}$ can be expressed in terms of its projection onto $\mathcal{A}(\bm{y})$ and its projection onto the orthogonal complement of $\mathcal{A}(\bm{y})$ as $\bm{x}=\bm{x}_{0}(\bm{y})+\bm{W}(\bm{y})\bm{v}+\bm{W}^{\perp}(\bm{y})\bm{u}$. Here, $\bm{W}^{\perp}(\bm{y})$ is a matrix with orthonormal columns that span $\text{Range}^{\perp}\{\bm{W}(\bm{y})\}$, and $\bm{u}$ corresponds to the coordinates within that subspace. Now, the sliced posterior distribution corresponds to evaluating the posterior at $\bm{u}=0$ in this decomposition. Namely, it is given by [^1] The function $f^{\text{sliced}}$ is not a density function as it is not normalized.
$f^{\text{sliced}}(\bm{v}|\bm{y})=p_{\mathbf{x}|\mathbf{y}}\left(\bm{x}_{0}(\bm{y})+\bm{W}(\bm{y})\bm{v}\middle|\bm{y}\right).$ (6)
@@
In contrast, the PDD at a point $\bm{v}$ corresponds to the integral of the posterior over all points whose projection onto $\mathcal{A}(\bm{y})$ has coordinates $\bm{v}$, or in other words, it is the integral over $\bm{u}$ in this decomposition,
$p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})$ $=\int p_{\mathbf{x}|\mathbf{y}}\left(\bm{x}_{0}(\bm{y})+\bm{W}(\bm{y})\bm{v}+\bm{W}^{\perp}(\bm{y})\bm{u}\middle|\bm{y}\right)d\bm{u}.$ (7)
@@
There exist settings in which $f^{\text{sliced}}$ equals $p_{\mathbf{v}|\mathbf{y}}$ (up to a normalization constant). This is the case, for example, when $\mathbf{u}$ and $\mathbf{v}$ are statistically independent given $\mathbf{y}$. However, in imaging inverse problems, this seems to rarely be the case. Indeed, the distribution of natural images is known to be concentrated near a low-dimensional manifold within the high-dimensional ambient space. Therefore, any low-dimensional subspace almost always misses the high-density regions of the distribution. On the other hand, the PDD provides meaningful information regardless of whether the slice passes through the high-density regions of the posterior or not. This is illustrated in Fig. 3
@@
Figure 4: Architecture overview. The degraded image $\bm{y}$ is first fed to a pre-selected subspace extractor that outputs an input-adaptive subspace $\mathcal{A}(\bm{y})=\{\bm{W}(\bm{y}),\bm{x}_{0}(\bm{y})\}$. Afterwards, the degraded image $\bm{y}$ and the extracted subspace are fed to a feature extractor that outputs a feature vector $h(\bm{y})$. The resulting $h(\bm{y})$ modulates a lightweight MLP that outputs $p(\bm{v}|\bm{y})$ for any query $\bm{v}$. The resulting projected distribution can then be navigated to visualize posterior uncertainty.
@@
Figure 5: Gaussian mixture denoising. (a) Prior distribution $p_{\mathbf{x}}{(\bm{x})}$, and a sample from the joint distribution $p_{\mathbf{x},\mathbf{y}}{(\bm{x},\bm{y})}$. (b) Full posterior distribution $p_{\mathbf{x}|\mathbf{y}}{(\bm{x}|\bm{y})}$, and the selected 1D subspace $\mathcal{A}(\bm{y})=\{\bm{x}_{0}(\bm{y}),\bm{W}(\bm{y})\}$. (c) Sliced $f^{\text{sliced}}(\bm{v}|\bm{y})$ (gray) vs projected $p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})$ (green) posterior distribution along the 1D subspace (dashed green line in (b)). (d) Comparison of the GT and the learned projected posterior $p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})$.
3.2 Choosing the subspace
Selecting an appropriate subspace is crucial for producing a meaningful PPD visualization. The ideal subspace should be either 1D or 2D to facilitate plotting. Moreover, the origin $\bm{x}_{0}(\bm{y})$ and directions $\bm{W}(\bm{y})$ should be chosen such that they disclose as much posterior variance as possible, hopefully in a semantically meaningful manner. A natural choice is to take the origin to be the minimum mean square error (MMSE) prediction, $\bm{x}_{0}(\bm{y})\approx\mathbb{E}[\mathbf{x}|\mathbf{y}=\bm{y}]$, and the directions $\bm{w}_{1}(\bm{y}),\dots,\bm{w}_{K}(\bm{y})$ to be the top PCs of the posterior. NPPC [49] outputs both $\bm{x}_{0}(\bm{y})$ and $\bm{w}_{1}(\bm{y}),\dots,\bm{w}_{K}(\bm{y})$ in a single forward pass, rendering it a suitable candidate for fast training and inference. In Fig. 3, we exemplify this choice for the task of image inpainting where moving about the origin $\bm{x}_{0}(\bm{y})$ along the selected direction $\bm{w}_{1}(\bm{y})$ changes digit identity.
3.3 Architecture
@@
Here we propose to learn the conditional distribution $p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})$ using a conditional EBM. Unlike normalizing flows, EBMs do not require specialized architectures, and are thus more flexible in their design. EBMs are usually trained with variants of contrastive divergence (CD) [22, 8, 6, 70] or noise contrastive estimation (NCE) [19], with various recent efforts to stabilize and improve training [18, 15, 76]
@@
As depicted in Fig. 4, the architecture we use has two parts: A (heavy) feature extractor with a classifier-style architecture and a lightweight small MLP with a few linear layers. The feature extractor receives the measurement $\bm{y}$ and the set of vectors defining the subspace $\mathcal{A}(\bm{y})$, and outputs a 1D feature vector $\bm{h}(\bm{y})$. This feature vector is then fed along with a queried point $\bm{v}$ into the second part that outputs $p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})$.
@@
This segregation of the architecture serves the purpose of lowering computational complexity. The rationale behind it is the following: For a given measurement $\bm{y}$ we would like to query $p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})$ for many values of $\bm{v}$. This is true both during training when running MCMC chains to produce contrastive samples, and during testing when evaluating the PPD on a grid of $\bm{v}$’s for plotting. Therefore, for a given measurement $\bm{y}$ and subspace $\mathcal{A}(\bm{y})$, this inner separation of the architecture allows us to evaluate the feature extractor only once while the lightweight MLP can be queried dozens of times incurring only a minor computational burden. The conditioning mechanism we adopt here is similar to the one employed for timestep encoding in diffusion models [23], with adaptive shifting and scaling of features reminiscent of AdaIN [25].
@@
To train our conditional EBM, we use a variant of CD [22]. In its basic form, CD fits a given parametric function $E(\bm{v};\bm{\theta})$ to the (unnormalized) negative log distribution function $-\log\tilde{p}_{\mathbf{v}}(\bm{v})$. In each training step, the MCMC process (e.g., Langevin dynamics) is applied to a batch of data samples $\bm{v}$ to produce contrastive samples $\bm{v}_{\text{neg}}$. The model parameters $\bm{\theta}$ are then updated as
$\bm{\theta}_{t+1}=\bm{\theta}_{t}-\nabla_{\bm{\theta}}[E(\bm{v};\bm{\theta})-E(\bm{v}_{\text{neg}};\bm{\theta})]\lvert_{\bm{\theta}_{t}}.$ (8)
@@
This process gradually increases the likelihoods that the model assigns to the samples in the training set and decreases the likelihoods that the model assigns to their contrastive counterparts, until convergence. At test time, the resulting model can output the probability density $p_{\mathbf{v}}{(\bm{v})}$ (up to an unknown normalization constant) for any queried $\bm{v}$.

### v2

@@ Abstract
In ill-posed inverse problems, it is commonly desirable to obtain insight into the full spectrum of plausible solutions, rather than extracting only a single reconstruction. Information about the plausible solutions and their likelihoods is encoded in the posterior distribution. However, for high-dimensional data, this distribution is challenging to visualize. In this work, we introduce a new approach for estimating and visualizing posteriors by employing energy-based models (EBMs) over low-dimensional subspaces. Specifically, we train a conditional EBM that receives an input measurement and a set of directions that span some low-dimensional subspace of solutions, and outputs the probability density function of the posterior within that space. We demonstrate the effectiveness of our method across a diverse range of datasets and image restoration problems, showcasing its strength in uncertainty quantification and visualization. As we show, our method outperforms a baseline that projects samples from a diffusion-based posterior sampler, while being orders of magnitude faster. Furthermore, it is more accurate than a baseline that assumes a Gaussian posterior. Code is available at https://github.com/yairomer/PPDE
A Introduction
Figure S1: Informed uncertainty visualization. Point estimation methods receive a distorted image and output only a single solution, e.g., the MMSE estimator (top). NPPC [48] complements MMSE estimators with input-adaptive uncertainty directions (principal components of the posterior) without modeling output likelihood (middle). Our method (bottom) learns the input-adaptive projected posterior distribution, facilitating a likelihood-informed uncertainty visualization.
@@ A Introduction
Recently, several works proposed to visualize posterior uncertainty by traversing along the principal components (PCs) of the posterior [48, 4]. These visualizations shed light on the main modes of variation along which the solution could vary given the input, and have thus been demonstrated to be natural for signals with strong correlations, like images. These methods are also mathematically appealing since one-dimensional statistics of the projected posterior along the PCs (like variances or quantiles) provide a tighter and more accurate description of the underlying uncertainty [4]. Nonetheless, methods in this category do not provide estimates for the likelihoods of the projected solutions within the low-dimensional subspace. This hinders the user’s ability to interpret and visualize the effective support of the projected posterior.
@@
In this work, we propose to model the projected posterior using a conditional energy-based model (EBM) that receives a degraded measurement and an affine subspace (parameterized by an origin point and a set of directions) and can output the projection of the posterior probability for any queried point within that subspace. This allows us to visualize the projected posterior within any low-dimensional space (e.g., the one spanned by the dominant posterior PCs), facilitating informed navigation of posterior uncertainty in both one and two dimensions (Fig. S1). Our method, which we coin projected posterior distribution estimation (PPDE), is a general technique for uncertainty visualization that is seamlessly transferable across tasks and datasets. We demonstrate it with the affine subspace that passes through the point prediction provided by an MSE-optimized model and spanned by the posterior PCs provided by the NPPC method [48] and . However, our method is not constrained to this specific choice and can wrap around any point estimate and an accompanying set of directions.
@@
As we illustrate, our proposed visualization approach cannot be achieved with methods that explicitly learn the high-dimensional posterior distribution, like conditional invertible models (e.g., SRFlow [37]) or score-based models that enable the calculation of the exact likelihood using the probability flow ODE [65, 66]. In particular, the naive approach of evaluating and plotting the posterior along a 2D or 1D slice through the high-dimensional distribution gives meaningless results, as it will almost surely miss the high-probability regions in space (see Fig. S3 and the explanation in Sec. C.1). By contrast, our approach corresponds to projecting the distribution onto the 2D or 1D slice (i.e., integrating over all other dimensions), an effect that is practically impossible to achieve with models that output the density in the high-dimensional space.
@@
We demonstrate the practical benefit of our method on multiple inverse problems in imaging. Our learned posteriors are quantitatively compared to a Gaussian approximation obtained from [48], as well as to a baseline that projects samples generated by a diffusion-based posterior sampler and applies kernel-density estimation (KDE) in the low-dimensional space. In both cases, we show a significant improvement in sample log-likelihood across tasks and datasets. This illustrates the benefit offered by our method compared to works that approximate the posterior with a Gaussian [14, 46, 43, 49] or attempt to rely on posterior samplers [4], setting forth a new approach to proper uncertainty visualization.
B Related Work
@@
C Method
Figure S2: Notations illustration. An image $\bm{x}$ is projected onto a measurement-adaptive affine subspace $\mathcal{A}(\bm{y})$, and represented by its projection coefficients $\bm{v}=[v_{1},v_{2}]^{\top}$. The restored image within the subspace is denoted by $\bm{x}_{\text{restored}}$.
C.1 Projecting the posterior distribution
Our goal is to visualize the uncertainty when predicting a signal $\bm{x}$ based on measurements $\bm{y}$. In the context of imaging, $\bm{y}$ often represents a degraded version of $\bm{x}$ (e.g., noisy, blurry). We assume that $\bm{x}$ and $\bm{y}$ are realizations of random vectors $\mathbf{x}$ and $\mathbf{y}$, respectively. The desired uncertainty we aim to explore is therefore encapsulated in the posterior distribution $p_{\mathbf{x}|\mathbf{y}}(\bm{x}|\bm{y})$. As mentioned earlier, this distribution is typically defined over a high-dimensional space, and therefore visualizing it requires projection onto a lower-dimensional manifold (e.g., 1D or 2D).
@@ C.1 Projecting the posterior distribution
$\left\{\bm{x}:\bm{x}=\bm{x}_{0}(\bm{y})+\bm{W}(\bm{y})\bm{v},\;\bm{v}\in\mathbb{R}^{K}\right\},$ (S1)
@@
$\bm{W}(\bm{y})=\begin{bmatrix}|&|&&|\\ \bm{w}_{1}(\bm{y})&\bm{w}_{2}(\bm{y})&\dots&\bm{w}_{K}(\bm{y})\\ |&|&&|\end{bmatrix}.$ (S2)
@@
$\mathcal{A}(\bm{y})=\left\{\bm{x}_{0}(\bm{y}),W(\bm{y})\right\}$ (S3)
and with slight abuse of terminology, refer to $\mathcal{A}(\bm{y})$ as our affine subspace (or just subspace). Our method can be used along with any method of producing a subspace $\mathcal{A}(\bm{y})$ for a given input $\bm{y}$. Here we choose $\bm{x}_{0}(\bm{y})$ to be a minimum MSE (MMSE) predictor, and use the NPPC method [48] for producing $\bm{W}(\bm{y})$ (see Sec. C.2).
For a given input $\bm{y}$, our goal is to visualize the posterior distribution of the coordinates of $\mathbf{x}$ projected onto the subspace $\mathcal{A}(\bm{y})$, which we denote by
$\mathbf{v}=\bm{W}(\bm{y})^{\top}(\mathbf{x}-\bm{x}_{0}(\bm{y})).$ (S4)
@@
$\bm{x}_{\text{restored}}=\bm{x}_{0}(\bm{y})+\bm{W}(\bm{y})\bm{v}.$ (S5)
Figure S2 summarizes our notations.
@@
Figure S3: Posterior slicing vs. projection. (a) Schematic illustrating the difference between the slice and the projection of a high-dimensional manifold onto a low-dimensional subspace. (b) Sliced and projected posterior comparison for a 1D affine subspace defined by the posterior mean and the first PC in the task of digit inpainting. The sliced posterior has been computed using an EBM.
@@
It is instructive to note the difference between the PPD and a sliced posterior distribution (i.e., the high-dimensional posterior evaluated on $\mathcal{A}(\bm{y})$). Any $\bm{x}$ can be expressed in terms of its projection onto $\mathcal{A}(\bm{y})$ and its projection onto the orthogonal complement of $\mathcal{A}(\bm{y})$ as $\bm{x}=\bm{x}_{0}(\bm{y})+\bm{W}(\bm{y})\bm{v}+\bm{W}^{\perp}(\bm{y})\bm{u}$. Here, $\bm{W}^{\perp}(\bm{y})$ is a matrix with orthonormal columns that span $\text{Range}^{\perp}\{\bm{W}(\bm{y})\}$, and $\bm{u}$ corresponds to the coordinates within that subspace. Now, the sliced posterior distribution corresponds to evaluating the posterior at $\bm{u}=0$ in this decomposition. Namely, it is given by [^1] The function $f^{\text{sliced}}$ is not a density function as it is not normalized.
$f^{\text{sliced}}(\bm{v}|\bm{y})=p_{\mathbf{x}|\mathbf{y}}\left(\bm{x}_{0}(\bm{y})+\bm{W}(\bm{y})\bm{v}\middle|\bm{y}\right).$ (S6)
@@
In contrast, the PPD at a point $\bm{v}$ corresponds to the integral of the posterior over all points whose projection onto $\mathcal{A}(\bm{y})$ has coordinates $\bm{v}$. In other words, the PPD is the marginal distribution of $\bm{v}$ produced by integrating out $\bm{u}$,
$p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})$ $=\int p_{\mathbf{x}|\mathbf{y}}\left(\bm{x}_{0}(\bm{y})+\bm{W}(\bm{y})\bm{v}+\bm{W}^{\perp}(\bm{y})\bm{u}\middle|\bm{y}\right)d\bm{u}.$ (S7)
@@
There exist settings in which $f^{\text{sliced}}$ equals $p_{\mathbf{v}|\mathbf{y}}$ (up to a normalization constant). This is the case, for example, when $\mathbf{u}$ and $\mathbf{v}$ are statistically independent given $\mathbf{y}$. However, in imaging inverse problems, this seems to rarely be the case. Indeed, the distribution of natural images is known to be concentrated near a low-dimensional manifold within the high-dimensional ambient space. Therefore, any low-dimensional subspace almost always misses the high-density regions of the distribution. On the other hand, the PPD provides meaningful information regardless of whether the slice passes through the high-density regions of the posterior or not. This is illustrated in Fig. S3.
@@
Figure S4: Architecture overview. The degraded image $\bm{y}$ is first fed to a pre-selected subspace extractor that outputs an input-adaptive subspace $\mathcal{A}(\bm{y})=\{\bm{W}(\bm{y}),\bm{x}_{0}(\bm{y})\}$. Afterward, the degraded image $\bm{y}$ and the extracted subspace are fed to a feature extractor that outputs a feature vector $h(\bm{y})$. The resulting $h(\bm{y})$ modulates a lightweight MLP that outputs $p(\bm{v}|\bm{y})$ for any query $\bm{v}$. The resulting projected distribution can then be navigated to visualize posterior uncertainty.
C.2 Choosing the subspace
Selecting an appropriate subspace is crucial for producing a meaningful PPD visualization. The ideal subspace should be either 1D or 2D to facilitate visualization. Moreover, the origin $\bm{x}_{0}(\bm{y})$ and directions $\bm{W}(\bm{y})$ should disclose as much posterior variance as possible, hopefully in a semantically meaningful manner. A natural choice is to take the origin to be the minimum mean square error (MMSE) prediction, $\bm{x}_{0}(\bm{y})\approx\mathbb{E}[\mathbf{x}|\mathbf{y}=\bm{y}]$, and the directions $\bm{w}_{1}(\bm{y}),\dots,\bm{w}_{K}(\bm{y})$ to be the top $K$ PCs of the posterior. NPPC [48] outputs both $\bm{x}_{0}(\bm{y})$ and $\bm{w}_{1}(\bm{y}),\dots,\bm{w}_{K}(\bm{y})$, rendering it a suitable candidate for fast training and inference. In Fig. S3, we exemplify this choice for the task of image inpainting where moving about the origin $\bm{x}_{0}(\bm{y})$ along the selected direction $\bm{w}_{1}(\bm{y})$ changes digit identity.
C.3 Architecture
@@
Figure S5: Gaussian mixture denoising. (a) Prior distribution $p_{\mathbf{x}}{(\bm{x})}$, and a sample from the joint distribution $p_{\mathbf{x},\mathbf{y}}{(\bm{x},\bm{y})}$. (b) Full posterior distribution $p_{\mathbf{x}|\mathbf{y}}{(\bm{x}|\bm{y})}$, and the selected 1D subspace $\mathcal{A}(\bm{y})=\{\bm{x}_{0}(\bm{y}),\bm{W}(\bm{y})\}$. (c) Sliced $f^{\text{sliced}}(\bm{v}|\bm{y})$ (purple) vs. projected $p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})$ (green) posterior distribution along the 1D subspace (dashed gray line in (b)). (d) Comparison of the GT and the learned projected posterior.
@@
We propose to learn the conditional distribution $p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})$ using a conditional EBM. Unlike normalizing flows, EBMs do not require specialized architectures, and are thus more flexible in their design. EBMs are usually trained with variants of contrastive divergence (CD) [22, 8, 6, 69] or noise contrastive estimation (NCE) [19], with various recent efforts to stabilize and improve training [18, 15, 76].
@@
As depicted in Fig. S4, we propose to design the architecture using two parts: A (heavy) feature extractor with a classifier-style architecture and a lightweight small MLP with a few linear layers. The feature extractor receives the measurement $\bm{y}$ and the set of vectors defining the subspace $\mathcal{A}(\bm{y})$, and outputs a 1D feature vector $\bm{h}(\bm{y})$. This feature vector is then fed along with a queried point $\bm{v}$ into the second part that outputs $p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})$. Using this architecture we expect the intermediate feature vector to encapsulate all the information regarding the 2D distribution and have the MLP just translate this data into a distribution function.
@@
This segregation of the architecture serves the purpose of lowering computational complexity. The rationale behind it is the following: For a given measurement $\bm{y}$, we would like to query $p_{\mathbf{v}|\mathbf{y}}(\bm{v}|\bm{y})$ for many values of $\bm{v}$. This is true both during training when running MCMC chains to produce contrastive samples, and during testing when evaluating the PPD on a grid of $\bm{v}$’s for plotting. Therefore, for a given measurement $\bm{y}$ and subspace $\mathcal{A}(\bm{y})$, this inner separation of the architecture allows us to evaluate the feature extractor only once while the lightweight MLP can be queried dozens of times incurring only a minor computational burden. The conditioning mechanism we adopt here is similar to the one employed for timestep encoding in diffusion models [23], with adaptive shifting and scaling of features reminiscent of AdaIN [25].
@@
To train our conditional EBM, we use a variant of CD [22]. In its basic form, CD fits a given parametric function $E(\bm{v};\bm{\theta})$ to the (unnormalized) negative log distribution function $-\log\tilde{p}_{\mathbf{v}}(\bm{v})$. In each training step, an MCMC process (e.g., Langevin dynamics) is applied to a batch of data samples $\bm{v}$ to produce contrastive samples $\bm{v}_{\text{neg}}$. The model’s parameters $\bm{\theta}$ are then updated as
$\bm{\theta}_{t+1}=\bm{\theta}_{t}-\nabla_{\bm{\theta}}\left.\left[\mathbb{E}_{\bm{v}}[E(\bm{v};\bm{\theta})]-\mathbb{E}_{\bm{v}_{\text{neg}}}[E(\bm{v}_{\text{neg}};\bm{\theta})]\right]\right|_{\bm{\theta}_{t}}.$ (S8)
@@
This process gradually increases the likelihood that the model assigns to the samples in the training set and decreases the likelihood that the model assigns to their contrastive counterparts, until convergence. At test time, the resulting model can output the probability density $p_{\mathbf{v}}{(\bm{v})}$ (up to an unknown normalization constant) for any queried $\bm{v}$.